1.    The following worksheet contains cost and revenue data for koyel Shoe Company:

 

Total 15,000 Pairs of Shoes

Per pair of                Shoes

Sales Revenue

Tk. 600,000

Tk. 40

Variable expenses:

 

 

     Invoice cost

202,500

Tk. 13.5

     Sales commission

67,500

      4.5

Total variable expenses

Tk. 270,000

Tk.18

Contribution margin

Tk. 330,000

Tk. 22

Fixed expenses:

 

 

     Advertising

Tk. 30,000

 

     Rent

20,000

 

     Salaries

100,000

 

Total fixed expenses

Tk. 150,000

 

Net Income

Tk. 180,000

 

Required:

(a)  Compute the company’s degree of operating leverage at the present level of sales.

(b)  Assume that through a more intense effort by the sales staff, the company’s sales increase

by 8% next year. By what percentage would you expect net operating income to increase? Use the degree of operating leverage to obtain your answer.

(c)  Verify your answer to (b) by preparing a new contribution format income statement showing an 8% increase in sales.



1
Expert's answer
2022-03-10T07:04:35-0500

Question (a). Compute the company’s degree of operating leverage at the present level of sales.


Solution (a)


To compute the company’s degree of operating leverage at the present level of sales, we use the formula


"Degree\,of\,Operating\,Leverage = \frac{{Contribution\,M\arg in}}{{Net\,Income}}\"


"Degree\,of\,Operating\,Leverage = \frac{{Tk 330,000}}{{Tk 180,000}}\"


"Degree\,of\,Operating\,Leverage =1.833333...."



Question (b). Assume that through a more intense effort by the sales staff, the company’s sales increase by 8% next year. By what percentage would you expect net operating income to increase? Use the degree of operating leverage to obtain your answer.


Solution (b)

"Operating\,Income\,Increase = Operating\,Leverage \times 8\% \"


"Operating\,Income\,Increase = 1.83333..... \times 8\% \"


"Operating\,Income\,Increase =14.66666....\% \"



Question (c). Verify your answer to (b) by preparing a new contribution format income statement showing an 8% increase in sales.


Solution (c)


Sales Revenue "= 15,000 \u00d7 Tk. 40 = Tk. 600,000" 

The variable expenses are

Invoice Cost "= 15,000 \u00d7 Tk. 13.5 = Tk. 202,500"

Sales Commission "= 15,000 \u00d7 Tk. 4.5 = Tk. 67,500"

Hence, total variable expenses are "= Tk. 202,500 + Tk. 67,500 = Tk. 270,000"

 

Contribution margin "= 15,000 \u00d7 Tk. 22 = Tk. 330,000"

 

Now Fixed Expenses are

Advertising "= Tk. 30,000"

Rent "= Tk. 20,000"

Salaries "= Tk. 100,000"

Total Fix Expenses

"= Tk 30,000 + Tk. 20,000 + Tk. 100,000\\ = Tk. 150,000"

 Net Income "= Tk. 180,000"